OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −(x + 6) y − 4 = 2(x + 6) y + 6 = −(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

@help123please.

OpenStudy (anonymous):

first we have to find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-6,4) x1 = -6 and y1 = 4 (2,0) x2 = 2 and y2 = 0 now lets sub.. m = (0 - 4) / (2 - (-6) m = -4/8 m = -1/2 <-- this is your slope now we will use the point slope formula : y - y1 = m(x - x1) using slope 1/2 and points (-6,4) y - 4 = -1/2(x - (-6) y - 4 = -1/2(x + 6) I am getting a different answer then your answer choices.